1. You walk up a staircase from floor 5 to floor 6 at constant velocity. Each floor is 4 m
above the other. Your mass is 70 kg.
a) How much WORK do you do?
b) Does it depend on how steep the stairs are, that is, is it different if there are more
steps each of which is less high?
c) What if you slide your shoes along the steps dramatically increasing your friction,
how does that change the WORK that you do?
d) If you slide your shoes along the steps dramatically increasing your friction how
does that change your GRAVITATIONAL POTENTIAL ENERGY when you arrive at floor 6?
Answer #1

a) work done = change in potential energy

= mg(h2 - h1)

= 70*9.8*4 = 2744 J

b) it only depends on the height , so design does not depend

c) work done would be higher as you have to work against friction

d) the gravitational potential energy would remain same as it depends on the height moved